# A method for sparse and robust independent component analysis

### v2

@@ Appendix B Proof of Theorem 4
To see that this holds, we first observe that the quantity inside the square root can be written as
$(1-\|v\|)^{2}+\|v\|\{2-\sqrt{2(1+2\varepsilon)}\},$
where $2(1+2\varepsilon)<4$, making (5) well-defined. Now, some simplification shows that (5) is equivalent to
$1+2u_{n}^{\prime}v+\|v\|^{2}<1+\|v\|^{2}-\sqrt{2(1+2\varepsilon)}\|v\|,$
which in turn is equivalent to $u_{n}^{\prime}v<-\|v\|(0.5+\varepsilon)^{1/2}$, which we next show to hold. Because $\|P_{n}v\|^{2}=(u_{n}^{\prime}v)^{2}$, our assumption writes as $(u_{n}^{\prime}v)^{2}>\|v\|^{2}(0.5+\varepsilon)$. As $\|u_{n}+v\|\leq 1$ implies that $u_{n}^{\prime}v\leq 0$, the case $u_{n}^{\prime}v>\|v\|(0.5+\varepsilon)^{1/2}$ cannot occur in the assumption and, consequently, it must hold that $u_{n}^{\prime}v<-\|v\|(0.5+\varepsilon)^{1/2}$, finally showing that (5) is true.
We next search still for a simpler lower bound for (5). Observe first that our three assumptions
$(u_{n}^{\prime}v)^{2}>\|v\|^{2}(0.5+\varepsilon),\quad\|u_{n}+v\|\leq 1,\quad u_{n}^{\prime}v\geq-1,$
imply that $1\leq u_{n}^{\prime}v\leq 0$ and, consequently, that $\|v\|<(0.5+\varepsilon)^{-1/2}$. We then claim that,
$1-\left\{1+\|v\|^{2}-\sqrt{2(1+2\varepsilon)}\|v\|\right\}^{1/2}\geq\|v\|(\sqrt{0.5+\varepsilon}-\sqrt{0.5-\varepsilon}).$
This inequality is equivalent to
@@
which is a quadratic in $\|v\|$ and can easily be verified to hold when $\|v\|<(0.5+\varepsilon)^{-1/2}$. Consequently in the sub-case (I), we have for $T_{n}$ the bound
$T_{n}\geq\rho_{n}\|v\|^{2}(\sqrt{0.5+\varepsilon}-\sqrt{0.5-\varepsilon})^{2}.$
What remains, is tackling the case where $\|P_{n}v\|^{2}/\|v\|^{2}>0.5+\varepsilon$ and $\|u_{n}+v\|>1$. In this case it holds that
$|\|u_{n}+v\|-1|\geq\sqrt{0.5+\varepsilon}\|v\|.$
To see this, we note that the inequality is equivalent to having
$2y\geq(C-1)x^{2}+2\sqrt{C}x,$
when $C=0.5+\varepsilon$ and $x,y\in\mathbb{R}$ satisfy $y^{2}>Cx^{2}$, $x\geq 0$, $2y>-x^{2}$ and $y\geq-1$.
Consequently, in the sub-case (II), we have the lower bound
$T_{n}\geq\rho_{n}(0.5+\varepsilon)\|v\|^{2}.$
Combining now all three lower bounds with the choice $\varepsilon=0.1$ and observing that $(\sqrt{0.6}-\sqrt{0.4})^{2}<0.6$ and that $\|A_{n}^{-1}\|_{2}^{-1}\|G_{n}^{-1/2}\|_{2}^{-1}\|m\|_{2}\leq\|A_{n}G_{n}^{1/2}m\|_{2}$, we get the desired claim.
@@
$\min\{C_{1n},C_{2n}\}\|A_{n}^{-1}\|_{2}^{-2}\|G_{n}^{-1/2}\|_{2}^{-2}$
$\geq$ $\min\left\{\frac{0.2\rho(\rho-\psi)}{3\psi},\frac{1}{2}\rho\right\}\frac{1}{18\rho^{2}}\Omega^{2}\alpha$
$=:$ $C,$
where $C>0$. Then, for $n\geq n_{0}$, if $\mathcal{A}_{n}$ holds, we have, by Lemma 2,
$f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})$
$\geq$ $\|m\|_{2}(C\|m\|_{2}-\varepsilon\sqrt{p}).$
Now this implies that any candidate minimizer $b_{n,0}+m$ of Version I must have $\|m\|_{2}<(2\varepsilon\sqrt{p})/C$. This is because, if $\|m\|_{2}\geq(2\varepsilon\sqrt{p})/C$, we have
$f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})\geq\frac{2\varepsilon^{2}p}{C}>0,$
showing that any such $b_{n,0}+m$ cannot be a minimizer of the objective function of Version I. Thus the result is proven. ∎
The consistency $b_{n,\lambda_{n}}-b_{n,0}\rightarrow_{p}0$ now follows directly from Lemma 3 and the earlier fact that $P(\mathcal{A}_{n})\rightarrow 1$ as $n\rightarrow\infty$.
@@
Assume that $\lambda_{n}\rightarrow 0$. Then we have,
$\|b_{n,\lambda_{n}}-b_{n,0}\|=o_{p}(1),$
as $n\rightarrow\infty$.
@@
Assume that $a_{n},\lambda_{n}\rightarrow 0$ be such that $\lambda_{n}/a_{n}\rightarrow 0$. Then we have,
$a_{n}^{-1}\|b_{n,\lambda_{n}}-b_{n,0}\|=o_{p}(1),$
as $n\rightarrow\infty$.
@@
$G_{n}(Ax_{1},\ldots,Ax_{n})$ $=AG_{n}(x_{1},\ldots,x_{n})A^{\prime},$
$M_{n}(Ax_{1},\ldots,Ax_{n})$ $=AM_{n}(x_{1},\ldots,x_{n})A^{\prime},$
@@
$c_{n}\|a_{n,0}-\alpha\|_{2}=\mathcal{O}_{p}\left(1\right)\quad\mbox{and}\quad c_{n}\|b_{n,0}-\beta\|_{2}=\mathcal{O}_{p}\left(1\right).$
@@
By the affine equivariance of $G,M$, the solution vector $b_{n,0}$ of Version II transforms as $b_{n,0}\mapsto(A^{-1})^{\prime}b_{n,0}$ under the transformation $(x_{1},\ldots,x_{n})\mapsto(Ax_{1},\ldots,Ax_{n})$ [55]. The equivalent also holds for the solution vector of Version III. Now, as
$\|b_{n,0}-\beta\|_{2}\leq\|A^{\prime}\|_{2}\|(A^{-1})^{\prime}b_{n,0}-(A^{-1})^{\prime}\beta\|_{2},$
where $\|\cdot\|_{2}$ for a matrix argument denotes the spectral norm, the result $\|b_{n,0}-\beta\|_{2}=\mathcal{O}_{p}(1/c_{n})$ follows once we show that $\|(A^{-1})^{\prime}b_{n,0}-(A^{-1})^{\prime}\beta\|_{2}=\mathcal{O}_{p}(1/c_{n})$ for any single choice of $A$.
Choose then $A=O^{\prime}G^{-1/2}$ where $O$ is any orthogonal matrix containing the eigenvectors of $G^{-1/2}MG^{-1/2}$ as its columns. Calling the diagonal matrix of respective eigenvalues $\Lambda$, we thus have $AGA^{\prime}=I_{p}$ and $AMA^{\prime}=\Lambda$. Consequently, we may in the following assume that $G=I_{p}$ and that $M$ is diagonal with strictly positive diagonal elements.
@@
we obtain, correcting the sign of $u_{n}$ if necessary, that $u_{n1}-1=\mathcal{O}_{p}(1/c_{n}^{2})$ and, consequently, that
$\|u_{n}-e_{1}\|_{2}=\mathcal{O}_{p}\left(\frac{1}{c_{n}}\right).$
By Corollary 1, we have $b_{n,0}=G_{n}^{-1/2}u_{n}$ and $\beta=e_{1}$. Arguing as in the proof of Lemma 1 in [57], we obtain $\|G_{n}^{-1/2}-I_{p}\|_{2}=\mathcal{O}_{p}(1/c_{n})$, finally giving us,
$\|b_{n,0}-\beta\|_{2}\leq$ $\|b_{n,0}-G_{n}^{-1/2}e_{1}\|_{2}+\|G_{n}^{-1/2}e_{1}-e_{1}\|_{2}$
$\leq$ $\|G_{n}^{-1/2}\|_{2}\|b_{n,0}-e_{1}\|_{2}+\|G_{n}^{-1/2}-I_{p}\|_{2}$
$=$ $\mathcal{O}_{p}(1/c_{n}).$
@@
Combining Corollary 3 and Lemma 4 via the triangle inequality, we obtain the following main result of this section.

### v3

@@ Appendix B Proof of Theorem 4
To see that this holds, we first observe that the quantity inside the square root can be written as $(1-\|v\|)^{2}+\|v\|\{2-\sqrt{2(1+2\varepsilon)}\}$, where $2(1+2\varepsilon)<4$, making (5) well-defined. Now, simplification shows that (5) is equivalent to $1+2u_{n}^{\top}v+\|v\|^{2}<1+\|v\|^{2}-\sqrt{2(1+2\varepsilon)}\|v\|$, which in turn is equivalent to $u_{n}^{\top}v<-\|v\|(0.5+\varepsilon)^{1/2}$, which we next show to hold. Because $\|P_{n}v\|^{2}=(u_{n}^{\top}v)^{2}$, our assumption writes as $(u_{n}^{\top}v)^{2}>\|v\|^{2}(0.5+\varepsilon)$. As $\|u_{n}+v\|\leq 1$ implies that $u_{n}^{\top}v\leq 0$, the case $u_{n}^{\top}v>\|v\|(0.5+\varepsilon)^{1/2}$ cannot occur in the assumption and, consequently, it must hold that $u_{n}^{\top}v<-\|v\|(0.5+\varepsilon)^{1/2}$, finally showing that (5) is true.
We next search still for a simpler lower bound for (5). Observe first that our three assumptions $(u_{n}^{\top}v)^{2}>\|v\|^{2}(0.5+\varepsilon)$, $\|u_{n}+v\|\leq 1$ and $u_{n}^{\top}v\geq-1$ imply that $1\leq u_{n}^{\top}v\leq 0$ and, consequently, that $\|v\|<(0.5+\varepsilon)^{-1/2}$. We then claim that the following inequality holds, $1-\left\{1+\|v\|^{2}-\sqrt{2(1+2\varepsilon)}\|v\|\right\}^{1/2}\geq\|v\|(\sqrt{0.5+\varepsilon}-\sqrt{0.5-\varepsilon})$. This inequality is equivalent to
@@
which is a quadratic in $\|v\|$ and can easily be verified to hold when $\|v\|<(0.5+\varepsilon)^{-1/2}$. Consequently in the sub-case (I), we have for $T_{n}$ the bound $T_{n}\geq\rho_{n}\|v\|^{2}(\sqrt{0.5+\varepsilon}-\sqrt{0.5-\varepsilon})^{2}$.
What remains, is tackling the case where $\|P_{n}v\|^{2}/\|v\|^{2}>0.5+\varepsilon$ and $\|u_{n}+v\|>1$. In this case it holds that $|\|u_{n}+v\|-1|\geq\sqrt{0.5+\varepsilon}\|v\|$. To see this, we note that the inequality is equivalent to having $2y\geq(C-1)x^{2}+2\sqrt{C}x$, when $C=0.5+\varepsilon$ and $x,y\in\mathbb{R}$ satisfy $y^{2}>Cx^{2}$, $x\geq 0$, $2y>-x^{2}$ and $y\geq-1$. Consequently, in the sub-case (II), we have the lower bound $T_{n}\geq\rho_{n}(0.5+\varepsilon)\|v\|^{2}$.
Combining now all three lower bounds with the choice $\varepsilon=0.1$ and observing that $(\sqrt{0.6}-\sqrt{0.4})^{2}<0.6$ and that $\|A_{n}^{-1}\|_{2}^{-1}\|G_{n}^{-1/2}\|_{2}^{-1}\|m\|\leq\|A_{n}G_{n}^{1/2}m\|$, we get the desired claim.
@@
$\min\{C_{1n},C_{2n}\}\|A_{n}^{-1}\|_{2}^{-2}\|G_{n}^{-1/2}\|_{2}^{-2}\geq\min\left\{\frac{0.2\rho(\rho-\psi)}{3\psi},\frac{1}{2}\rho\right\}\frac{1}{18\rho^{2}}\Omega^{2}\alpha=:C,$
where $C>0$. Then, for $n\geq n_{0}$, if $\mathcal{A}_{n}$ holds, we have, by Lemma 2, that $f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})\geq\|m\|(C\|m\|-\varepsilon\sqrt{p})$. Now this implies that any candidate minimizer $b_{n,0}+m$ of Version I must have $\|m\|<(2\varepsilon\sqrt{p})/C$. This is because, if $\|m\|\geq(2\varepsilon\sqrt{p})/C$, we have $f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})\geq 2\varepsilon^{2}p/C>0$, showing that any such $b_{n,0}+m$ is not a minimizer of the objective function of Version I. Thus the result is proven. ∎
The consistency $b_{n,\lambda_{n}}-b_{n,0}\rightarrow_{p}0$ now follows from Lemma 3 and the earlier fact that $P(\mathcal{A}_{n})\rightarrow 1$ as $n\rightarrow\infty$.
@@
Assume that $\lambda_{n}\rightarrow 0$. Then we have $\|b_{n,\lambda_{n}}-b_{n,0}\|=o_{p}(1)$, as $n\rightarrow\infty$.
@@
Assume that $a_{n},\lambda_{n}\rightarrow 0$ be such that $\lambda_{n}/a_{n}\rightarrow 0$. Then we have $a_{n}^{-1}\|b_{n,\lambda_{n}}-b_{n,0}\|=o_{p}(1)$, as $n\rightarrow\infty$.
@@
$G_{n}(Ax_{1},\ldots,Ax_{n})=AG_{n}(x_{1},\ldots,x_{n})A^{\top}\quad\mbox{and}\quad M_{n}(Ax_{1},\ldots,Ax_{n})=AM_{n}(x_{1},\ldots,x_{n})A^{\top},$
@@
$c_{n}\|a_{n,0}-\alpha\|=\mathcal{O}_{p}\left(1\right)\quad\mbox{and}\quad c_{n}\|b_{n,0}-\beta\|=\mathcal{O}_{p}\left(1\right).$
@@
By the affine equivariance of $G,M$, the solution vector $b_{n,0}$ of Version II transforms as $b_{n,0}\mapsto(A^{-1})^{\top}b_{n,0}$ under the transformation $(x_{1},\ldots,x_{n})\mapsto(Ax_{1},\ldots,Ax_{n})$ [56]. The equivalent also holds for the solution vector of Version III. Now, as $\|b_{n,0}-\beta\|\leq\|A^{\top}\|_{2}\|(A^{-1})^{\top}b_{n,0}-(A^{-1})^{\top}\beta\|$, where we recall that $\|\cdot\|_{2}$ denotes the spectral norm, the result $\|b_{n,0}-\beta\|=\mathcal{O}_{p}(1/c_{n})$ follows once we show that $\|(A^{-1})^{\top}b_{n,0}-(A^{-1})^{\top}\beta\|=\mathcal{O}_{p}(1/c_{n})$ for any single choice of $A$.
Choose then $A=O^{\top}G^{-1/2}$ where $O$ is any orthogonal matrix containing the eigenvectors of $G^{-1/2}MG^{-1/2}$ as its columns. Calling the diagonal matrix of respective eigenvalues $\Lambda$, we thus have $AGA^{\top}=I_{p}$ and $AMA^{\top}=\Lambda$. Consequently, we may in the following assume that $G=I_{p}$ and that $M$ is diagonal with strictly positive diagonal elements.
@@
we obtain, correcting the sign of $u_{n}$ if necessary, that $u_{n1}-1=\mathcal{O}_{p}(1/c_{n}^{2})$ and, consequently, that $\|u_{n}-e_{1}\|=\mathcal{O}_{p}(1/c_{n})$. By Corollary 1, we have $b_{n,0}=G_{n}^{-1/2}u_{n}$ and $\beta=e_{1}$. Arguing as in the proof of Lemma 1 in [58], we obtain $\|G_{n}^{-1/2}-I_{p}\|_{2}=\mathcal{O}_{p}(1/c_{n})$, finally giving us,
$\|b_{n,0}-\beta\|\leq$ $\|b_{n,0}-G_{n}^{-1/2}e_{1}\|+\|G_{n}^{-1/2}e_{1}-e_{1}\|\leq\|G_{n}^{-1/2}\|_{2}\|b_{n,0}-e_{1}\|+\|G_{n}^{-1/2}-I_{p}\|_{2}=\mathcal{O}_{p}(1/c_{n}).$
@@
Combining Corollary 3 and Lemma 4 via the triangle inequality, we obtain the main result of this section.